Jason borrows $20,000 at an effective annual rate of 13.00% and promises to pay it back over 11 quarters with equal quarterly installments. Out of the sixth repayment (i.e. the repayment at t=6), how many dollars of the repayment go toward paying the interest portion and how many dollars go toward paying the principal portion?
Answer #1

Loan Amount EAR Nper 11 20,000 AMORTIZATION SCHEDULE 13% Period Initial Balance Repayment Interest Principal Ending Balance 2

Formulae

20000 0.13 Period 1618 Loan Amount 1619 EAR 1620 Nper 1621 1622 1623 1624 1625 1626 1627 =C1620+1 =C1621+1 =C1622+1 =C1623+1